# Wannier states and spin supersolid physics in the triangular antiferromagnet K2Co(SeO3)2

### v1

@@ III Model and theory
In the study of real magnetic systems, we encounter two key challenges: extracting a model that accurately describes the most relevant aspects of the system’s physics, and then solving it. This task is “ouroboric” in nature, as determining the model’s parameters requires solving it first and comparing the results with experimental observations. Since most models cannot be solved exactly, this makes the process highly non-trivial. A reasonable approach is to propose a minimal model that captures the essential physical features of the system, at least within a certain energy scale, and solve it using well-justified approximations. Following this methodology, and building on the results outlined in the introduction and in our previous work [28], we propose modeling the magnetic properties of K2Co(SeO3)2 using an XXZ Hamiltonian,
@@
where $h=g\mu_{B}\mu_{0}H$, $\mu_{0}H$ is the external magnetic field, $\mu_{B}$ is the Bohr magneton and $g=7.9$ the Landé g-factor. The exchange interactions determined in [28] through high-field magnetometry measurements were further refined by detailed comparison with theoretical calculations, as discussed in the following section. This process yielded updated values of $J_{zz}=3.1\,\text{meV}$, $J_{xy}=0.217\,\text{meV}$, and an anisotropy ratio of $\alpha=J_{xy}/J_{zz}=0.07$. The refinement was essential to achieve a precise agreement with the specific heat and magnetization curves presented in subsequent sections.
Let us first highlight the classical and semi-classical predictions of this model and evaluate how well they align with experiments. In the strictly Ising limit ($J_{xy}=0$) and in the absence of the external magnetic field ($\mu_{0}H=0$), the Hamiltonian becomes a triangular lattice Ising model, exhibiting an extensively degenerate ground state manifold [21]. The ground state subspace is spanned by configurations where no triangular plaquette has all three spins aligned in parallel. In the introduction, we referred to this subspace as the “Wannier space” $\mathcal{W}$. An infinitesimal magnetic field applied along the $z$-direction lifts the degeneracy and favors three possible “uud” ground states, where two spin sublattices align parallel to the field, and the third sublattice has spins oriented antiparallel to the field. This implies that the magnetization changes discontinuously from zero to one-third of the saturation value.
As we move away from the Ising limit $J_{xy}\ll J_{zz}$ in the absence of the external field, the ground state manifold of the classical spin ($S\to\infty$) Hamiltonian is two-dimensional. The first dimension emerges from the continuous U(1) symmetry associated with global spin rotations around the $z$-axis. The second dimension arises from an accidental continuous degeneracy, which corresponds to the freedom in selecting one of the three polar angles for the spins within the magnetic unit cell [44, 45]. These angles, measured with respect to the $z$-axis as illustrated in Figure 1(b), satisfy the following relationships,
@@
with $\vartheta_{A}$ arbitrarily chosen. The application of an infinitesimal magnetic field along the $z$-direction lifts the accidental degeneracy, selecting the Y-shaped ground state configuration depicted in Figure 1(a).
@@
Although the classical ground state of the Hamiltonian in Eq. 1 qualitatively matches the state of K2Co(SeO3)2, it completely fails to predict the zero-field magnetization. The classical solution yields a magnetization of
$\frac{M}{M_{\rm sat}}=\frac{1}{3}\frac{(J_{zz}-J_{xy})}{(J_{zz}+J_{xy})}\ \xrightarrow{\alpha\ll 1}\frac{1}{3}$ (3)
@@
which strongly contrasts with the zero magnetization observed experimentally, as shown in Figure 3. The absence of ground-state magnetization can be attributed to a significant reduction in the magnitude of the two quasi-parallel spins, approximately by a factor of two. However, such a pronounced reduction—resulting in the ordered moment being only 50% of its full value—indicates that quantum corrections are not adequately captured by a perturbative $1/S$ expansion. Although we cannot entirely rule out the possibility of antiferromagnetic interlayer correlations characterized by $J_{\perp}$, the absence of a jump or noticeable change in the slope of the magnetization curve suggests that each individual layer has zero net magnetization.
@@
Figure 3: Magnetization of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ as a function of applied field at $T=0.19~{}\text{K}$. Circles are experimental data from Ref. [28]. The solid curve is a guide for the eye. Red squares are QMC results (this work). The dashed line represents one-third of the experimental saturation magnetization.
Kleine et al. [44] demonstrated that quantum fluctuations lift the classical accidental ground state degeneracy by stabilizing the coplanar Y-order. This ordering pattern breaks two symmetries simultaneously: it spontaneously breaks the continuous U(1) symmetry of global spin rotations by selecting a specific orientation in the $xy$-plane, and it breaks the discrete ${\text{C}}_{3}$ lattice symmetry (corresponding to $32\pi/$ 3 rotations) by designating one of the three sublattices to host the minority spin. Notably, in the absence of a magnetic field, there is an additional breaking of time-reversal ${\text{Z}}_{2}$ symmetry, corresponding to the up or down orientation of the minority spin. It is interesting to explore the consequences of these different symmetry breakings for the expected thermodynamic phase transitions in the pure 2D model.
In view of the distinct energy scales governing the stabilization of Ising (solid) and XY (BEC) orderings, it is clear that the solid or Ising ordering, characterized by the spontaneous breaking of ${\text{C}}_{3}\otimes Z_{2}$ symmetry at zero field and ${\text{C}}_{3}$ symmetry at finite field, should occur at a higher critical temperature. At zero field, we expect this solid ordering to emerge through a sequence of two Berezinskii-Kosterlitz-Thouless (BKT) transitions, with $T_{\rm BKT,1}>T_{\rm BKT,2}$. A critical phase exists for $T_{\rm BKT,2}<T<T_{\rm BKT,1}$, while a massive phase with long-range Ising order develops for $T<T_{\rm BKT,2}$ [46].
The behavior changes significantly in the presence of a finite magnetic field, where only the ${\text{C}}_{3}$ symmetry is broken. Under these conditions, the critical phase acquires long-range Ising order and merges with the massive phase below $T_{\rm BKT,2}$. The two BKT transitions are replaced by a single second-order phase transition in the 2D Potts universality class at temperature $T_{\rm c1}$. As the magnetic field approaches zero, $T_{\rm c1}$ converges to $T_{\rm BKT,1}$. Lastly, the XY order parameter is expected to undergo a BKT transition at a much lower temperature, $T_{\rm BKT,3}\ll T_{\rm BKT,2}$. When finite inter-layer coupling is included, the transition evolves into a three-dimensional XY phase transition, resulting in a phase with long-range XY order. Consequently, K2Co(SeO3)2 is expected to exhibit weak thermodynamic signatures of these phase transitions at zero field, particularly in measurements of specific heat ($C_{v}$). Figure 4(a) presents the zero-field specific heat of K2Co(SeO3)2 measured up to high temperatures (open circles). The phonon contribution is approximated by the measured specific heat of the isostructural nonmagnetic counterpart K2Mg(SeO3)2 (solid squares). An empirical scale factor of 1.08 was applied in order to match the heat capacities of the two systems in the temperature range 50--100 K. The magnetic heat capacity $C_{\text{mag}}$ of K2Co(SeO3)2 is plotted as a function of temperature in Figure 4(b). Here the phonon contribution and a nuclear Schottky feature at the lowest temperatures were subtracted from the total heat capacity. Signatures of three BKT transitions are observed (dashed lines), which are well aligned with the theoretical predictions.
Figure 4: (a) Heat capacity measured in K2Co(SeO3)2 (shaded circles) and in the non-magnetic counterpart K2Mg(SeO3)2 (solid squares) as a function of temperature. (b) Temperature dependence of the magnetic contribution to specific heat in K2Co(SeO3)2. The phonon contribution was estimated based on the K2Mg(SeO3)2 data. A 1.08 scale factor was applied to the latter, to match the heat capacities of the two materials in the range 50 to 100 K. The two prominent peaks correspond to the two higher-temperature BKT transitions. A weaker feature at low temperature marks the third transition. Note the logarithmic temperature scale. Actual BKT transition temperatures are typically lower than the corresponding specific heat maximum [47].
While thermal fluctuations can be qualitatively captured by a classical spin model, the same does not hold true for quantum fluctuations. Using a simple spin-wave calculation, Kleine observed a substantial reduction in the total ground-state magnetization, nearly driving it to zero. While this result aligns closely with expectations for the quantum $S=1/2$ limit, it also raises concerns about the validity of the $1/S$ expansion. Specifically, the first quantum correction to the dipole moment renormalization is nearly as large as the full moment, highlighting the limitations of this approach in effectively addressing the problem.
@@
At low energies, $\omega\sim J_{xy}$, the spin-wave calculation predicts two degenerate magnon bands: one corresponding to a true Goldstone mode associated with the broken U(1) symmetry, and the other a pseudo-Goldstone mode arising from the accidental continuous degeneracy of the classical ground state [see solid lines in Figure 5]. A third, nearly dispersionless band (not shown in the Figure) appears at higher energies, $\omega\sim J_{zz}$. In the $S=1/2$ limit, this mode corresponds to excitations into states in the subspace ${\cal W}^{\perp}$, which is orthogonal to the Wannier subspace ${\cal W}$.
Higher-order $1/S$ corrections are expected to gap out the pseudo-Goldstone mode [48]. This is in qualitative agreement with our ultra-high-resolution inelastic neutron scattering measurement at the AMATERAS spectrometer using an incident neutron energy $E_{i}$ = 1.69 meV (more discussions in Sec. V), where a tiny energy gap of 60 $\mu$eV is detected at the K point [Figure 5 inset]. However, the comparison of the linear spin wave theory (LSWT) dispersion with the INS data in Figure 5 and the strong reduction of the ordered moment suggests that the $1/S$ expansion is insufficient to accurately capture the dispersion and spectral weight distribution of shorter-wavelength excitations. Consequently, we will adopt an alternative approach to account for the non-perturbative effects of strong quantum fluctuations.
Figure 5: False-color plot of inelastic neutron scattering intensity measured in K2Co(SeO3)2 in zero applied field and $T=0.07$ K as a function of energy and wave vector transfers along high-symmetry directions. The incident neutron energy is $E_{i}$ = 1.69 meV. The energy resolution is 23 $\mu$eV at the elastic line. This is to be compared to the 36 $\mu$eV resolution of the experiments described in Ref. [28]. The data are integrated over ±$\pm 0.5$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Fig. 20. No background has been subtracted. Both longitudinal and transverse spin fluctuations contribute to scattering, the intensity being proportional to $S^{zz}(\mathbf{q},\omega)+S^{\bot\bot}(\mathbf{q},\omega)$. White solid lines are the dispersion curves calculated by linear spin wave theory. The white hexagon represents the Brillouin zone boundary. $\Gamma$, K and M label high symmetry points in the reciprocal space. Inset: a zoom-in view of the low-energy spectrum near the K point with a different color scale to highlight the gapped pseudo-Goldstone excitation. Labels on the top axis indicate the position of energy cuts shown in Figure 11.
Since our approach is based on a perturbative expansion in a different small parameter, $J_{xy}/J_{zz}$, it is essential to first verify that the excitation spectrum of K2Co(SeO3)2 exhibits two well-separated energy scales. This separation should be evident in the measured temperature dependence of the entropy. As calculated by Wannier, the zero-point entropy density of the pure Ising model is approximately $\Delta S/R\approx 0.323$, which is about half of $\ln{2}$ )—the saturation entropy for a spin-$1/2$ system.
@@
At zero field and temperature, quantum fluctuations induced by first-order corrections in $J_{xy}$ remove the extensive ground state degeneracy by selecting the above discussed ground state sub-manifold within the Wannier subspace. The Wannier entropy should then be recovered at a crossover temperature $T=T_{\rm cr}$ of order $J_{xy}/k_{B}$ and the entropy curve
@@
should exhibit a plateau at this temperature. Experimentally, $\Delta S(T)$ of K2Co(SeO3)2 is obtained by integrating from the lowest temperature measured, as plotted against temperature in Figure 6. A plateau is observed, which confirms this separation of energy scales and reveals the value of the crossover temperature $T_{\rm cr}\simeq 2.5~{}\text{K}$. The relatively narrow temperature range of the observed plateau can be attributed to the moderate separation of energy scales, with $J_{xy}$ being only ten times smaller than $J_{zz}$.
Figure 6: Temperature dependence of magnetic entropy in $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ in zero applied field, as obtained by integrating the measured specific heat and subtracting the phonon contribution determined from experiments on $\text{K}_{2}\text{Mg}(\text{Se}\text{O}_{3})_{2}$. The dashed lines indicate the Wannier entropy and the expected high-temperature value $\ln{2}$ ). The inset provides a zoomed-in view to highlight the plateau at the Wannier entropy. The 10% discrepancy between the high-temperature data and the expected saturation value is attributed to imperfect phonon subtraction. This effect is expected to be much smaller at temperatures below 10 K.
By exploiting this separation of energy scales, we derive a low-energy effective model ${\cal H}_{\rm eff}$ through degenerate perturbation theory to first order in $J_{xy}$. Our approach is based on the observations initially made by A. Sen et al. [49] for the non-frustrated case and subsequently generalized by Fa Wang and collaborators [50]. The derivation begins by projecting the original Hamiltonian in Eq. (1) onto the Wannier subspace ${\cal W}$ to produce an effective low-energy model. Wang et al. demonstrate the existence of a unitary transformation that reverses the sign of the $xy$ term of the effective low-energy model while preserving both the eigenvalue spectrum and the diagonal correlation functions in the $S^{z}$ basis. This transformation yields a Hamiltonian free from the sign problem, enabling the use of QMC simulations to compute both static and dynamic properties. In the following section, we detail this unitary transformation.

### v2

@@ II.4 Hamiltonian model
In the study of real magnetic systems, we encounter two key challenges: extracting a model that accurately describes the most relevant aspects of the system’s physics, and then solving it. This task is “ouroboric” in nature, as determining the model’s parameters requires solving it first and comparing the results with experimental observations. Since most models cannot be solved exactly, this makes the process highly non-trivial. A reasonable approach is to propose a minimal model that captures the essential physical features of the system, at least within a certain energy scale, and solve it using well-justified approximations. Following this methodology, and building on the results outlined in the previous sections and in our previous work [33], we propose modeling the magnetic properties of K2Co(SeO3)2 using an XXZ Hamiltonian,
@@
where $h=g_{zz}\mu_{B}\mu_{0}H$, $\mu_{0}H$ is the external magnetic field, $\mu_{B}$ is the Bohr magneton and $g_{zz}=7.9$ the ${z}$-component of the Landé g-tensor. The exchange interactions determined in [33] through high-field magnetometry measurements were further refined by detailed comparison with theoretical calculations, as discussed below. This process yielded updated values of $J_{zz}=3.1\,\text{meV}$, $J_{xy}=0.217\,\text{meV}$, and an anisotropy ratio of $\alpha=J_{xy}/J_{zz}=0.07$. The refinement was essential to achieve a precise agreement with the specific heat and magnetization curves presented in subsequent sections.
We begin by examining the zero-temperature classical and semi-classical predictions for the Hamiltonian model in Eq. 1 and assessing their agreement with experimental observations. Given the small ratio $\alpha$, it is natural to first consider the strictly Ising limit ($J_{xy}=0$). In the absence of an external magnetic field ($H=0$), the Hamiltonian reduces to an antiferromagnetic triangular Ising model, which exhibits an extensively degenerate ground-state manifold [22]. This manifold is spanned by configurations in which no triangular plaquette has all three spins aligned in parallel. As introduced earlier, we refer to this subspace as the “Wannier space” $\mathcal{W}$.
An infinitesimal magnetic field applied along the $z$-direction lifts this degeneracy, stabilizing three distinct “uud” ground states, where two spin sublattices align parallel to the field while the third sublattice orients antiparallel to it. Consequently, the magnetization undergoes a discontinuous jump from zero to one-third of the saturation value.
As we move away from the Ising limit ($J_{xy}\ll J_{zz}$) and in the absence of an external field, the ground-state manifold of the classical spin ($S\to\infty$) Hamiltonian becomes two-dimensional. The first dimension emerges from the continuous U(1) Hamiltonian symmetry associated with global spin rotations around the $z$-axis. The second dimension arises from an accidental continuous degeneracy, which corresponds to the freedom in selecting one of the three polar angles for the spins within the magnetic unit cell [44, 45, 46, 32]. These angles, measured relative to the $z$-axis, as shown in Figure 1(b), satisfy the following relationships:
@@
with $\vartheta_{A}$ arbitrarily chosen. Eventually, the application of an infinitesimal magnetic field along the $z$-direction lifts the accidental degeneracy, selecting the Y-shaped ground state configuration depicted in Figure 1(a).
@@
Although the classical ground state of the Hamiltonian in Eq. 1 qualitatively matches the state of K2Co(SeO3)2, it completely fails to predict the magnetization at zero field. The classical solution gives a magnetization of
$\frac{M(H=0)}{M_{\rm sat}}=\frac{1}{3}~{}\frac{1-\alpha}{1+\alpha}\ \xrightarrow{\alpha\ll 1}\frac{1}{3}$ (3)
@@
which stands in stark contrast to the zero magnetization observed experimentally, as shown in Figure 12. Although we cannot entirely rule out the possibility of antiferromagnetic interlayer correlations characterized by a coupling constant $J_{\perp}$, the absence of a jump or noticeable change in the slope of the magnetization curve suggests that each individual layer has zero net magnetization.
@@
Figure 12: Field dependence of magnetization in K2Co(SeO3)2. Magnetization of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ as a function of applied field at $T=0.19~{}\text{K}$. Circles are experimental data from Ref. [33]. The solid curve is a guide for the eye. Red squares are QMC results (this work). The blue curve is the LSWT calculation. The dashed line represents one-third of the experimental saturation magnetization.
Moving now to a semi-classical approach, spin-wave calculations by Kleine et al. [44] demonstrated that quantum fluctuations lift the classical accidental ground-state degeneracy, stabilizing the coplanar Y-order. Additionally, they observed a significant reduction in the total ground-state magnetization, nearly driving it to zero, which aligns closely with our experimental results. However, while the zero-field magnetization predicted by LSWT is close to our experimental findings, the magnetization curve $M(\mu_{0}H)$ exhibits a significant discrepancy with the experimental data [see Figure 12], particularly in predicting a lower critical field for the transition to the 1/3 plateau.
@@
At low energies ($\omega\sim J_{xy}$) and zero field, spin-wave calculations predict two degenerate magnon bands. One corresponds to a Goldstone mode associated with the broken U(1) symmetry, arising from the selection of a specific orientation for the spins in the $xy$-plane. The other is a “pseudo-Goldstone” mode, resulting from the accidental continuous degeneracy of the classical ground state [see solid lines in Figure 5]. Higher-order $1/S$ corrections are expected to gap out the pseudo-Goldstone mode [47, 48]. This is in qualitative agreement with our ultra-high-resolution inelastic neutron scattering measurement, where a tiny energy gap of 60 $\mu$eV is detected at the K point [Figure 5 inset]. At higher energies ($\omega\sim J_{zz}$), a third, nearly dispersionless band emerges [Figure 7]. This mode corresponds to excitations into states within the subspace $\mathcal{W}^{\perp}$, which is orthogonal to the Wannier subspace $\mathcal{W}$, as illustrated in Figure 8.
While LSWT accurately describes the excitation velocity near the K-point, it fails to capture the local minima in the dispersion around the M-point observed in the INS data [Figure 5]. Moreover, the significant reduction in the ordered moment found by Kleine suggests that a $1/S$ expansion is inadequate for accurately describing the dispersion and spectral weight distribution of shorter-wavelength excitations. In particular, the first quantum correction to the dipole moment renormalization is nearly as large as the full moment, underscoring the absence of a suitable expansion parameter in this approach.
It is now insightful to examine the thermodynamics, i.e., the $T>0$ regime. As calculated by Wannier, the pure Ising model at zero field exhibits a macroscopic residual entropy (zero-temperature entropy) of $\Delta S/R\approx 0.323$, approximately half of $\ln 2$—the full saturation entropy of a spin-$\frac{1}{2}$ system.
@@
In our model, quantum fluctuations arising from first-order corrections in $J_{xy}$ are expected to lift the extensive ground-state degeneracy by selecting a specific ground-state submanifold within the Wannier subspace. Consequently, the Wannier entropy should be recovered at a crossover temperature $T_{\rm cr}\sim J_{xy}/k_{B}$, where the entropy curve
@@
should display a plateau. This behavior is precisely observed in K2Co(SeO3)2, as shown in Figure 4. The relatively narrow temperature range of the plateau can be attributed to the moderate separation of energy scales, with $J_{xy}$ being “only” an order of magnitude smaller than $J_{zz}$.
As demonstrated in the seminal work of José et al. [49], identifying the different spontaneously broken symmetries is essential for understanding phase transitions in spin models. Given the distinct energy scales governing Ising (solid) and XY (BEC) orderings, the former occurs at a higher critical temperature.
At zero field, the Ising ordering breaks the discrete $\text{C}_{3}$ lattice symmetry (associated with $32\pi/$ 3 rotations) by selecting one of the three sublattices to host the minority spin. It also breaks time-reversal ($\mathbb{Z}_{2}$) symmetry, as the minority spin can point either up or down [see Figure 13]. Consequently, this ordering emerges through two Berezinskii-Kosterlitz-Thouless (BKT) transitions at $T_{\text{BKT},1}$ and $T_{\text{BKT},2}$, with $T_{\text{BKT},1}>T_{\text{BKT},2}$. A critical phase exists for $T_{\text{BKT},2}<T<T_{\text{BKT},1}$, while below $T_{\text{BKT},2}$, a massive phase with long-range Ising order sets in. This behavior is evident in our high-temperature specific heat data [Figure 3(b)].
At even lower temperatures, the XY order—associated with U(1) spin symmetry breaking—undergoes a separate BKT transition at $T_{\text{BKT},3}\ll T_{\text{BKT},2}$. In our experimental data, this transition appears as a subtle elbow in the inset of Figure 3(b).
Since only the $\text{C}_{3}$ symmetry breaks for finite magnetic field, the critical phase acquires long-range Ising order and merges with the massive phase below $T_{\text{BKT},2}$. Simultaneously, the two BKT transitions are replaced by a single second-order phase transition in the 2D Potts universality class at $T_{\text{c}1}$, which converges to $T_{\text{BKT},1}$ as the field vanishes.
Although real systems, which are inherently three-dimensional, are expected to exhibit a single phase transition at $T_{c1}\simeq T_{{\rm BKT},1}$ replacing the upper two transitions, and a three-dimensional XY phase transition at $T_{c2}\simeq T_{{\rm BKT},3}$ replacing the lower BKT transition, the anomaly in the specific heat associated with these transitions is too subtle to be detected in our experiments. Consequently, K2Co(SeO3)2 approximates extremely well the expected phase transition of the 2D XXZ model (Eq. 1) and shows qualitative consistency with the predictions of previous studies [46, 50, 45, 51].
Figure 13: Discrete lattice symmetries. Sketch of the discrete symmetries broken during the emergence of Ising (solid) order at zero field (large box) and at non-zero field (small boxes). The choice between the top or bottom small box depends on the orientation of the external field.
Up to this point, we have presented arguments, supported by both the current literature and our experimental data, demonstrating that K2Co(SeO3)2 is well described by the XXZ Hamiltonian (Eq. 1). Specifically, we highlighted that LSWT accurately captures the renormalization of the ordered moment at zero temperature and zero field but fails to reproduce the $M(\mu_{0}H)$ curve. Additionally, while thermal fluctuations are qualitatively well described by a classical spin model, our INS data reveal that quantum fluctuations significantly impact the excitation spectrum, causing deviations from the semi-classical prediction. This suggests the need for a fully quantum approach to properly capture the system’s behavior. Although a $1/S$ expansion could be pursued, its convergence is uncertain due to the lack of a well-defined small expansion parameter. In the remainder of this work, we will take a different approach to solving the full quantum Hamiltonian.
II.5 Low energy theory
Exploiting the clear energy separation discussed earlier, explicitly reflected in the ratio $\alpha$, we treat this parameter as a perturbative one. This approach follows the initial observations of A. Sen et al. [52] for the non-frustrated case ($J_{xy}<0$), later independently generalized by Fa Wang[29] and H. C. Jiang [30] to the frustrated regime.
To capture the low-energy physics, we construct an effective Hamiltonian, ${\cal H}_{\text{eff}}$, by projecting the original Hamiltonian (Eq. (1)) onto the Wannier subspace ${\cal W}$. Wang et al. demonstrate the existence of a unitary transformation that reverses the sign of the $J_{xy}$ term in ${\cal H}_{\text{eff}}$ while preserving both the eigenvalue spectrum and diagonal correlations in the $S^{z}$ basis. This transformation eliminates the QMC sign problem, allowing us to harness this powerful numerical technique to investigate both static and dynamic properties. In the following, we provide a detailed discussion of this transformation.